OpenStudy (anonymous):

What rule changes the input numbers to output numbers? input / output 1/-6 2/-4 3/-2 4/0

OpenStudy (ash2326):

Let o/p variable be y and i/p be x let y=mx+c m and c are constants , we have to find when i/p is 4 o/p is 0 y=0 , x=4 0=4m+c so c=-4m when x=3 y=-2 -2=3m+c c=-4m -2=3m-4m -2=-m m=2 we have c=-4m c=-8 so we get y=2x-8

OpenStudy (anonymous):

thank you soooo much

OpenStudy (ash2326):

welcome, did you understand it?

OpenStudy (anonymous):

i think so